### v1

@@ 1 Introduction
Protein fitness optimization seeks to improve the functionality of a protein by altering its amino acid sequence, to achieve a desired biological property of interest called “fitness” – for instance, stability, binding affinity, or catalytic efficiency. It requires searching through a vast combinatorial space (referred to as the “fitness landscape”), where the number of possible sequences grows exponentially with the sequence length $d$, while only a small subset of these sequences exhibit meaningful biological functionality Hermes et al. (1990). Traditionally, protein fitness optimization has been addressed through directed evolution Romero & Arnold (2009), mimicking natural evolution in the laboratory with a time-consuming, yet narrow random exploration of the fitness landscape. This highlights the need for efficient in-silico methods capable of exploring the space of potential sequences, so as to prioritize promising candidates for experimental validation.
Many different approaches exist, from generative models Jain et al. (2022); Gruver et al. (2024); Notin et al. (2022), evolutionary greedy algorithms Sinai et al. (2020) to gradient-based sampling strategies, such as Gibbs With Gradients (GWG) Grathwohl et al. (2021); Emami et al. (2023) and smoothed GWG variants Kirjner et al. (2023). In general, these methods are challenged by the high-dimensional, sparse and discrete nature of the fitness landscape, characterized by ruggedness Van Cleve & Weissman (2015) due to epistasis as well as holes due to proteins with very low fitness Johnston et al. (2023); Sinai & Kelsic (2020). To deal with the inherently discrete nature of amino acids, computational methods tend to rely on discretized processes, and thus must cope with issues such as non-smooth gradients and the vast diversity of possible sequences Kirjner et al. (2023); Frey et al. (2023).
@@
To construct a prior distribution over latent protein sequences, we leverage flow matching Lipman et al. (2023); Liu et al. (2022), a powerful generative modeling scheme that learns smooth, continuous representations amenable to gradient-based sampling and optimization. The prior is then integrated into a variational framework, making it possible to guide the search towards regions of high fitness with a fitness predictor in the form of a neural network, trained on a limited set of sequence mutations with associated fitness labels. See Figure 1. The versatility of the variational framework means that it can easily be tuned to different protein optimization tasks by suitably adapting the prior and the guidance function. We validate VLGPO on two public benchmarks for protein fitness optimization in limited data regimes, namely Adeno-Associated Virus (AAV) Bryant et al. (2021) and Green Fluorescent Protein (GFP) Sarkisyan et al. (2016), as suggested by Kirjner et al. (2023).
@@
  • We demonstrate state-of-the-art performance on established benchmarks for protein fitness optimization, namely AAV and GFP, targetting in particular tasks of medium and high difficulty in a limited data regime.
@@ 2 Related Work
We consider in-silico methodologies for protein fitness optimization at the sequence level, while emphasizing that there is an extensive stream of research centered around active learning frameworks, which iteratively integrate computational predictions with experimental validations Yang et al. (2025); Lee et al. (2024). While these methods are often considered the gold standard due to their integration of experimental feedback, our contribution is confined to purely computational strategies, thereby complementing existing active learning approaches.
@@
In a typical directed evolution setup Romero & Arnold (2009), biologists simulate nature in the laboratory by running multiple rounds of mutagenesis, searching through the local landscape by sampling a few mutations away from the current position in the sequence, effectively discovering new sequences through random walks in the sequence space. This process can be resource-intensive, as many mutations do not enhance the fitness of the starting sequence, and the vast number of combinations makes exhaustive exploration impractical. Therefore, many techniques have been introduced to address these challenges by employing surrogate models to guide the search more efficiently Sinai et al. (2020); Brookes et al. (2019); Trabucco et al. (2021); Ren et al. (2022). For instance, Adalead Sinai et al. (2020) uses a black-box predictor model to inform a greedy algorithm to prioritize mutations that are more likely to improve protein fitness.
@@
Moreover, generative methods have been explored for sequence generation. For instance, GFlowNets (GFN-AL) Jain et al. (2022) are designed to sample discrete objects, such as biological sequences, with probabilities proportional to a given reward function, facilitating the generation of diverse and high-quality candidates. Further, discrete diffusion models have been used for sequence sampling, where gradients in the hidden states of the denoising network help to guide the sequence design Gruver et al. (2024). Alternatively, a walk-jump algorithm was proposed to learn the distribution of one-hot encoded sequences using a single noise level and a Tweedie step to recover the sampled sequences after MCMC sampling Frey et al. (2023), which was further extended using gradient guidance in the noisy manifold Ikram et al. (2024). Likewise, also GWG proposes a strategy to obtain gradients for MCMC sampling of discrete sequences Grathwohl et al. (2021); Emami et al. (2023). The method Gibbs sampling with Graph-based Smoothing (GGS) builds upon this by additionally regularizing the noisy fitness landscape with graph-based smoothing Kirjner et al. (2023).
@@
Since the search space of protein sequences grows with sequence length, many works have started to recognize the significance of a latent space Praljak et al. (2023) or embedding spaces that are used in large-scale protein language models Lin et al. (2023). LatentDE Tran et al. (2024) for instance combines directed evolution with a smoothed space which allows to employ gradient ascent for protein sequence design in the latent space guided by a fitness predictor.
@@ 3.1 Protein Optimization
Therefore, we work with paired data $\mathcal{S}=\{(x_{i},y_{i})\}_{i=1}^{N}$ (note that $\mathcal{S}$ is only a subset of the entire data $\mathcal{S}^{*}$, see Table 1) and we use the parameterized convolutional neural network (CNN) $g:\mathcal{V}^{d}\to\mathbb{R}$ Kirjner et al. (2023) to infer the fitness for a given sequence. Specifically, we employ $g_{\phi}$ and $g_{\tilde{\phi}}$ as predictors, which are trained on a small subset of the data (see Section 4.1) without and with graph-based smoothing, respectively. Additionally, we use $g_{\psi}$ as an in-silico oracle for the final evaluation, trained on the entire paired data set $\mathcal{S}^{*}$. Note that all models share an identical architecture, differing only in their respective weights which we re-use without further training.
@@ 3.2 Generative Modeling
A recent class of generative models, known as diffusion models Ho et al. (2020); Song et al. (2021); Sohl-Dickstein et al. (2015), has shown remarkable success in generating high-quality data across various domains. These models work by progressively transforming simple noise distributions into complex data distributions through a series of iterative steps. During training, noise is systematically added to the data sample $x$ at varying levels, simulating a degradation process over time $t$. The model $\epsilon_{\theta}$ is then tasked with learning to reverse this process by predicting the added noise $\epsilon$ for $x_{t}$ at each step, effectively reconstructing the original data from the noisy observations. In detail, a model is trained to predict the added noise $\epsilon$ at each step by minimizing the objective
@@
Flow Matching. A more recent approach to generative modeling is given by the versatile framework of flow matching Lipman et al. (2023); Liu et al. (2022). Rather than removing noise from data samples, flow matching aims to model the velocity of the probability flow $\Psi_{t}$, which governs the dynamics of how one probability distribution is transformed into another over time. By learning the velocity field $u_{t}$ of the probability flow, the model $v_{\theta,t}$ captures the evolution of a simple base distribution at $t=0$ into a more complex target distribution $\mathrm{p}(x)$ at $t=1$, directly modeling the flow between them. Since the velocity field $u_{t}$ is intractable, it was shown in Lipman et al. (2023) that we can equivalently minimize the conditional flow matching loss:
@@ 3.3 Classifier guidance
Diffusion and flow-based methods can be guided by the log-likelihood gradient of an auxiliary classifier during generation, which enables conditional sampling Dhariwal & Nichol (2021). In detail, our goal is to sample from a conditional distribution $\mathrm{p}(x|g_{\phi}(x)=y)$, where $g_{\phi}$ represents the predictor and $y$ denotes the desired fitness value. We therefore adopt the framework of classifier guidance which allows for a decomposition of $\mathrm{p}(x|y)\sim\mathrm{p}(y|x)\mathrm{p}(x)$. Thus, we guide the sampling process by using the gradient of the predictor $g_{\phi}(x)$ with respect to the input sequence $x$. By introducing this gradient into the generative process, we can bias the sampling trajectory towards regions of the distribution that are more likely to yield sequences with the desired fitness value. The velocity field $v_{\theta,t}$ in the generative framework is modified to incorporate this guidance, yielding the following variational update:
@@
where $\nabla_{x}\log\mathrm{p}(y|x)\sim-\frac{1}{2}\nabla_{x}\|g_{\phi}(x)-y\|^{2}$ represents the gradient of the log-likelihood of a sequence $x$ having desired fitness $y$ and $\alpha_{t}$ is a scheduler dependent constant Zheng et al. (2023). For the classifier guidance we use the trained CNN-based predictors $g_{\phi}$ and the smoothed $g_{\tilde{\phi}}$ from Kirjner et al. (2023). Note that for guiding the process towards the highest fitness, $y$ is simply set to 1, which represents the highest fitness in the normalized fitness spectrum.
@@ 3.4 Latent space representation
So far, we introduced a general framework that in theory allows for sampling from a learned distribution of protein sequences. However, these sequences represent proteins that are combinations of a discrete set of amino acids. As a result, the underlying distribution of these sequences is likely sparse and complex, making it difficult to approximate or directly sample from in its original form. To overcome this limitation, we operate in an embedded latent space, where protein sequences are encoded as continuous representations. This is achieved by a VAE framework Kingma & Welling (2022), which maps discrete sequences to a continuous latent space through an encoder $\mathcal{E}:\mathcal{V}^{d}\mapsto\mathbb{R}^{l}$ and reconstructs them back to the original sequence space using a decoder network $\mathcal{D}:\mathbb{R}^{l}\mapsto\mathbb{R}^{d\times|\mathcal{V}|}$, where $l<<d$. Because of the discrete nature of amino acid tokens, the decoder produces logits, which are then mapped to tokens via an argmax operation. The training objective of the employed $\beta$-VAE Higgins et al. (2017) is given by the weighted Evidence Lower Bound (ELBO):
@@ 3.5 Variational Latent Generative Protein Optimization
From here on we will use the building blocks introduced earlier to describe our proposed VLGPO approach. We start by using our respective sequence data $x\sim\mathrm{p}(x)$ to train a VAE with encoder $\mathcal{E}$ and decoder $\mathcal{D}$ that compresses the higher-dimensional discrete protein sequences into a continuous latent space representation $z\sim\mathrm{p}(z|x)$. In order to model and sample from the learned latent space in an effective way, we train a flow matching model $v_{\theta,t}$ to learn the probability flow dynamics in the latent space. This network captures the transformation between a simple base distribution (e.g., Gaussian noise) at $t=0$ and the complex latent distribution of protein sequences at $t=1$. By integrating the learned flow from $t\in[0,1]$, we can efficiently generate new latent representations $z_{1}$, which can subsequently be decoded back into protein sequences via the VAE decoder, resulting in $x\sim\mathrm{p}_{\nu}(x|z_{1})=\mathcal{D}(z_{1})$, similar to Esser et al. (2024).
@@
Sampling from the posterior. Instead of merely sampling from the sequence distribution $\mathrm{p}(x)$, we also seek to generate high-fitness sequences. To achieve this, we condition the sampling process on a fitness score $y$, ideally set to the maximum value $y=1$, which is estimated by a predictor $\hat{y}=g_{\phi}(x)$. Although incorporating the likelihood term into the gradient, as shown in Equation 3, may appear straightforward, it is actually challenging because it becomes intractable due to the time-dependent nature of the diffusion/flow model. Moreover, explicitly including this term can push the generated samples off the current manifold related to time point $t$ Chung et al. (2022). Inspired by Chung et al. (2023); Ben-Hamu et al. (2024), we employ a scheme that evaluates the likelihood at $\hat{x}_{1}$, while the gradient of the likelihood is calculated at $x_{t}$, which has the effect of constraining the update to the same manifold. Therefore, at each of the $K$ sampling steps the model estimates $\hat{z}_{1}$ (Line 7, Algorithm 1) which is decoded to $\hat{x}_{1}$, the denoised version of the current decoded sample $x_{t}$, using the learned flow model $v_{\theta,t}$. The likelihood is then evaluated at $\hat{x}_{1}$ to reflect the target data distribution. However, the gradient of the likelihood, which guides the generative process, is computed with respect to $x_{t}$, leading to backpropagation through $v_{\theta,t}$. The hyperparameters $J$ and $\alpha_{t}$ denote the number of gradient descent steps on the likelihood and the guidance strength, respectively. Furthermore, as the predictor $g_{\phi}$ was trained in sequence space, our likelihood function changes to $\nabla_{x}\log\mathrm{p}(y|x_{1})\sim\frac{1}{2}\nabla_{z_{t}^{\prime}}\|g_{\phi}(\mathcal{D}(\hat{z}_{1}))-y\|^{2}$, i.e., we additionally have to backpropagate through the decoder. The pseudocode of our VLGPO can be found in Algorithm 1, a detailed visualization of the sampling scheme is reported in Figure 2.

### v2

@@ Abstract
protein design, variational methods, generative modeling
Figure 1: Overview of VLGPO sampling. The central section illustrates the VAE framework, showcasing protein sequences, their latent representations $z$, and the approximate posterior distribution. While the upper section depicts unconditional sampling from the prior $\mathrm{p}(x)$ using flow matching in the latent space, the lower section illustrates the modifications introduced by VLGPO during sampling. We additionally incorporate a likelihood term $\mathrm{p}(y|x)$ to condition on the fitness $y$, enabling sequence generation from the posterior distribution $\mathrm{p}(x|y)$ and facilitating sampling from high-fitness regions (as shown by the shifted and reshaped distribution).
@@ 1 Introduction
Protein fitness optimization seeks to improve the functionality of a protein by altering its amino acid sequence, to achieve a desired biological property of interest called “fitness” – for instance, stability, binding affinity, or catalytic efficiency. It requires searching through a vast combinatorial space (referred to as the “fitness landscape”), where the number of possible sequences grows exponentially with the sequence length $d$, while only a small subset of these sequences exhibit meaningful biological functionality (Hermes et al., 1990). Traditionally, protein fitness optimization has been addressed through directed evolution (Romero & Arnold, 2009), mimicking natural evolution in the laboratory with a time-consuming, yet narrow random exploration of the fitness landscape. This highlights the need for efficient in-silico methods capable of exploring the space of potential sequences, so as to prioritize promising candidates for experimental validation.
Many different approaches exist, from generative models (Jain et al., 2022; Gruver et al., 2024; Notin et al., 2022), evolutionary greedy algorithms (Sinai et al., 2020) to gradient-based sampling strategies, such as Gibbs With Gradients (GWG) (Grathwohl et al., 2021; Emami et al., 2023) and smoothed GWG variants (Kirjner et al., 2023). In general, these methods are challenged by the high-dimensional, sparse and discrete nature of the fitness landscape, characterized by ruggedness (Van Cleve & Weissman, 2015) due to epistasis as well as holes due to proteins with very low fitness (Johnston et al., 2023; Sinai & Kelsic, 2020). To deal with the inherently discrete nature of amino acids, computational methods tend to rely on discretized processes, and thus must cope with issues such as non-smooth gradients and the vast diversity of possible sequences (Kirjner et al., 2023; Frey et al., 2023).
@@
To construct a prior distribution over latent protein sequences, we leverage flow matching (Lipman et al., 2023; Liu et al., 2022), a powerful generative modeling scheme that learns smooth, continuous representations amenable to gradient-based sampling and optimization. The prior is then integrated into a variational framework, making it possible to guide the search towards regions of high fitness with a fitness predictor in the form of a neural network, trained on a limited set of sequence mutations with associated fitness labels. See Figure 1. The versatility of the variational framework means that it can easily be tuned to different protein optimization tasks by suitably adapting the prior and the guidance function. We validate VLGPO on two public benchmarks for protein fitness optimization in limited data regimes, namely Adeno-Associated Virus (AAV) (Bryant et al., 2021) and Green Fluorescent Protein (GFP) (Sarkisyan et al., 2016), as suggested by (Kirjner et al., 2023).
@@
  • We demonstrate state-of-the-art performance on established benchmarks for protein fitness optimization, namely AAV and GFP, targetting in particular tasks of medium and high difficulty in a limited data regime [^1] Source code available at
    https://github.com/uzh-dqbm-cmi/VLGPO.
@@ 2 Related Work
We consider in-silico methodologies for protein fitness optimization at the sequence level, while emphasizing that there is an extensive stream of research centered around active learning frameworks, which iteratively integrate computational predictions with experimental validations (Yang et al., 2025; Lee et al., 2024). While these methods are often considered the gold standard due to their integration of experimental feedback, our contribution is confined to purely computational strategies, thereby complementing existing active learning approaches.
@@
In a typical directed evolution setup (Romero & Arnold, 2009), biologists simulate nature in the laboratory by running multiple rounds of mutagenesis, searching through the local landscape by sampling a few mutations away from the current position in the sequence, effectively discovering new sequences through random walks in the sequence space. This process can be resource-intensive, as many mutations do not enhance the fitness of the starting sequence, and the vast number of combinations makes exhaustive exploration impractical. Therefore, many techniques have been introduced to address these challenges by employing surrogate models to guide the search more efficiently (Sinai et al., 2020; Brookes et al., 2019; Trabucco et al., 2021; Ren et al., 2022). For instance, Adalead (Sinai et al., 2020) uses a black-box predictor model to inform a greedy algorithm to prioritize mutations that are more likely to improve protein fitness.
@@
Moreover, generative methods have been explored for sequence generation. For instance, GFlowNets (GFN-AL) (Jain et al., 2022) are designed to sample discrete objects, such as biological sequences, with probabilities proportional to a given reward function, facilitating the generation of diverse and high-quality candidates. Further, discrete diffusion models have been used for sequence sampling, where gradients in the hidden states of the denoising network help to guide the sequence design (Gruver et al., 2024). Alternatively, a walk-jump algorithm was proposed to learn the distribution of one-hot encoded sequences using a single noise level and a Tweedie step to recover the sampled sequences after MCMC sampling (Frey et al., 2023), which was further extended using gradient guidance in the noisy manifold (Ikram et al., 2024). Likewise, also GWG proposes a strategy to obtain gradients for MCMC sampling of discrete sequences (Grathwohl et al., 2021; Emami et al., 2023). The method Gibbs sampling with Graph-based Smoothing (GGS) builds upon this by additionally regularizing the noisy fitness landscape with graph-based smoothing (Kirjner et al., 2023).
@@
Since the search space of protein sequences grows with sequence length, many works have started to recognize the significance of a latent space (Praljak et al., 2023) or embedding spaces that are used in large-scale protein language models (Lin et al., 2023). LatentDE (Tran et al., 2024) for instance combines directed evolution with a smoothed space which allows to employ gradient ascent for protein sequence design in the latent space guided by a fitness predictor.
@@ 3.1 Protein Optimization
Therefore, we work with paired data $\mathcal{S}=\{(x_{i},y_{i})\}_{i=1}^{N}$ (note that $\mathcal{S}$ is only a subset of the entire data $\mathcal{S}^{*}$, see Table 2) and we use the parameterized convolutional neural network (CNN) $g:\mathcal{V}^{d}\to\mathbb{R}$ (Kirjner et al., 2023) to infer the fitness for a given sequence. Specifically, we employ $g_{\phi}$ and $g_{\tilde{\phi}}$ as predictors, which are trained on a small subset of the data (see Section 4.1) without and with graph-based smoothing, respectively. Additionally, we use $g_{\psi}$ as an in-silico oracle for the final evaluation, trained on the entire paired data set $\mathcal{S}^{*}$. Note that all models share an identical architecture, differing only in their respective weights which we re-use without further training.
@@ 3.2 Generative Modeling
A recent class of generative models, known as diffusion models (Ho et al., 2020; Song et al., 2021; Sohl-Dickstein et al., 2015), has shown remarkable success in generating high-quality data across various domains. These models work by progressively transforming simple noise distributions into complex data distributions through a series of iterative steps. During training, noise is systematically added to the data sample $x$ at varying levels, simulating a degradation process over time $t$. The model $\epsilon_{\theta}$ is then tasked with learning to reverse this process by predicting the added noise $\epsilon$ for $x_{t}$ at each step, effectively reconstructing the original data from the noisy observations. In detail, a model is trained to predict the added noise $\epsilon$ at each step by minimizing the objective
@@
Flow Matching. A more recent approach to generative modeling is given by the versatile framework of flow matching (Lipman et al., 2023; Liu et al., 2022). Rather than removing noise from data samples, flow matching aims to model the velocity of the probability flow $\Psi_{t}$, which governs the dynamics of how one probability distribution is transformed into another over time. By learning the velocity field $u_{t}$ of the probability flow, the model $v_{\theta,t}$ captures the evolution of a simple base distribution at $t=0$ into a more complex target distribution $\mathrm{p}(x)$ at $t=1$, directly modeling the flow between them. Since the velocity field $u_{t}$ is intractable, it was shown in (Lipman et al., 2023) that we can equivalently minimize the conditional flow matching loss:
@@ 3.3 Classifier guidance
Diffusion and flow-based methods can be guided by the log-likelihood gradient of an auxiliary classifier during generation, which enables conditional sampling (Dhariwal & Nichol, 2021). In detail, our goal is to sample from a conditional distribution $\mathrm{p}(x|g_{\phi}(x)=y)$, where $g_{\phi}$ represents the predictor and $y$ denotes the desired fitness value. We therefore adopt the framework of classifier guidance which allows for a decomposition of $\mathrm{p}(x|y)\sim\mathrm{p}(y|x)\mathrm{p}(x)$. Thus, we guide the sampling process by using the gradient of the predictor $g_{\phi}(x)$ with respect to the input sequence $x$. By introducing this gradient into the generative process, we can bias the sampling trajectory towards regions of the distribution that are more likely to yield sequences with the desired fitness value. The velocity field $v_{\theta,t}$ in the generative framework is modified to incorporate this guidance, yielding the following variational update:
@@
where $\nabla_{x}\log\mathrm{p}(y|x)\sim-\frac{1}{2}\nabla_{x}\|g_{\phi}(x)-y\|^{2}$ represents the gradient of the log-likelihood of a sequence $x$ having desired fitness $y$ and $\alpha_{t}$ is a scheduler dependent constant (Zheng et al., 2023). Given the goal to maximize the fitness value of generated sequences, one could also set the gradient of the log-likelihood to $\frac{1}{2}\nabla_{x}\|g_{\phi}(x)\|^{2}$, or a similar suitable form. Both approaches perform very similar in practice. In this work, we chose the first version to demonstrate the possibility of steering sequences toward specific fitness values.
For the classifier guidance we use the trained CNN-based predictors $g_{\phi}$ and the smoothed $g_{\tilde{\phi}}$ from (Kirjner et al., 2023). Note that for guiding the process towards the highest fitness, $y$ is simply set to 1, which represents the highest fitness in the normalized fitness spectrum.
@@ 3.4 Latent space representation
So far, we introduced a general framework that in theory allows for sampling from a learned distribution of protein sequences. However, these sequences represent proteins that are combinations of a discrete set of amino acids. As a result, the underlying distribution of these sequences is likely sparse and complex, making it difficult to approximate or directly sample from in its original form. To overcome this limitation, we operate in an embedded latent space, where protein sequences are encoded as continuous representations. This is achieved by a VAE framework (Kingma & Welling, 2022), which maps discrete sequences to a continuous latent space through an encoder $\mathcal{E}:\mathcal{V}^{d}\mapsto\mathbb{R}^{l}$ and reconstructs them back to the original sequence space using a decoder network $\mathcal{D}:\mathbb{R}^{l}\mapsto\mathbb{R}^{d\times|\mathcal{V}|}$, where $l<<d$. Because of the discrete nature of amino acid tokens, the decoder produces logits, which are then mapped to tokens via an argmax operation. The training objective of the employed $\beta$-VAE (Higgins et al., 2017) is given by the weighted Evidence Lower Bound (ELBO):
@@
Note that while the VAE is trained with variational inference, VLGPO goes further by introducing additional generative modeling components: flow matching as described in Section 3.2 and classifier-guided sampling in Section 3.3. We show how these extensions refine the generation and help to better control it.
@@ 3.5 Variational Latent Generative Protein Optimization
From here on we will use the building blocks introduced earlier to describe our proposed VLGPO approach.
@@
We start by using our respective sequence data $x\sim\mathrm{p}(x)$ to train a VAE with encoder $\mathcal{E}$ and decoder $\mathcal{D}$ that compresses the higher-dimensional discrete protein sequences into a continuous latent space representation $z\sim\mathrm{p}(z|x)$. In order to model and sample from the learned latent space in an effective way, we train a flow matching model $v_{\theta,t}$ to learn the probability flow dynamics in the latent space. This network captures the transformation between a simple base distribution (e.g., Gaussian noise) at $t=0$ and the complex latent distribution of protein sequences at $t=1$. By integrating the learned flow from $t\in[0,1]$, we can efficiently generate new latent representations $z_{1}$, which can subsequently be decoded back into protein sequences via the VAE decoder, resulting in $x\sim\mathrm{p}_{\nu}(x|z_{1})=\mathcal{D}(z_{1})$, similar to (Esser et al., 2024).
Sampling from the posterior. Instead of merely sampling from the sequence distribution $\mathrm{p}(x)$, we also seek to generate high-fitness sequences. To achieve this, we condition the sampling process on a fitness score $y$, ideally set to the maximum value $y=1$, which is estimated by a predictor $\hat{y}=g_{\phi}(x)$.
@@
Although incorporating the likelihood term into the gradient, as shown in Equation 3, may appear straightforward, it is actually challenging because it becomes intractable due to the time-dependent nature of the diffusion/flow model. Moreover, explicitly including this term can push the generated samples off the current manifold related to time point $t$ (Chung et al., 2022). Inspired by (Chung et al., 2023; Ben-Hamu et al., 2024), we employ a scheme that evaluates the likelihood at $\hat{x}_{1}$, while the gradient of the likelihood is calculated at $x_{t}$, which has the effect of constraining the update to the same manifold. Therefore, at each of the $K$ sampling steps the model estimates $\hat{z}_{1}$ (Line 7, Algorithm 1) which is decoded to $\hat{x}_{1}$, the denoised version of the current decoded sample $x_{t}$, using the learned flow model $v_{\theta,t}$. The likelihood is then evaluated at $\hat{x}_{1}$ to reflect the target data distribution. However, the gradient of the likelihood, which guides the generative process, is computed with respect to $x_{t}$, leading to backpropagation through $v_{\theta,t}$. Furthermore, as the predictor $g_{\phi}$ was trained in sequence space, our likelihood function changes to
$\nabla_{x}\log\mathrm{p}(y|x_{1})\sim\frac{1}{2}\nabla_{z_{t}^{\prime}}\|g_{\phi}(\mathcal{D}(\hat{z}_{1}))-y\|^{2},$ (5)
where we additionally have to backpropagate through the decoder. The pseudocode of our VLGPO can be found in Algorithm 1, a detailed visualization of the sampling scheme is reported in Figure 2. The hyperparameters $J$ and $\alpha_{t}$ denote the number of gradient descent steps on the likelihood and the guidance strength, respectively.